Two numbers are in the ratio 2:5. If 6 is subtracted from their sum, the result is 50. What is the larger number?
Answer by sachi(548)

Two numbers are in the ratio 2:5.
so let the nos be 2x & 5x
If 6 is subtracted from their sum, the result is 50.
so 2x+5x-6=50
or 7x=50+6=56
or x=8
the larger number is 5*8=40
ans